OpenStudy (anonymous):

In how many ways can a teacher arrange 10 students in the front row if there are 60 total students?

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (michele_laino):

I think that the requested way is given by the number of all possible subsets of 10 students that we can make using 60 students

OpenStudy (michele_laino):

do you know what is that number?

OpenStudy (anonymous):

its permutations

OpenStudy (michele_laino):

no, since we have to make subsets of 10 students, using 60 students in total

OpenStudy (anonymous):

oh

OpenStudy (michele_laino):

do you know the binomial coefficients?

OpenStudy (anonymous):

no

OpenStudy (michele_laino):

the number which I mentioned above is the subsequent binomial coefficient: \[\left( {\begin{array}{*{20}{c}} {60} \\ {10} \end{array}} \right) = \frac{{60!}}{{10!\left( {60 - 10} \right)!}}\]

OpenStudy (anonymous):

so the answer is 6.974076e+139 @Michele_Laino

OpenStudy (michele_laino):

I got this: 75394027566 since we have: \[\left( {\begin{array}{*{20}{c}} {60} \\ {10} \end{array}} \right) = \frac{{60!}}{{10!\left( {60 - 10} \right)!}} = \frac{{60!}}{{10!\;50!}} = {\text{75394027566}}\]